# Semi-definite optimization of the measured relative entropies of quantum states and channels

### v2

@@ 1.1 Background
The relative entropy [KL51] and its generalization to Rényi relative entropy [R6́1] are important distinguishability measures in information theory, finding direct operational meaning in hypothesis testing tasks [Che52, Ste, Che56, Str62, Hoe65, HK89] while being used to construct other entropic measures like mutual information and conditional entropy [Sib69, Aug78, Csi95]. There are a number of quantum generalizations of these quantities [Ume62, Pet85, Pet86, MLDS+13, WWY14], finding operational meaning in quantum hypothesis testing tasks [HP91, ON00, Nag06, ACM+07, Hay07, ANSV08, NS09] while being used to construct other entropic measures like quantum mutual information and conditional entropy (see, e.g., [KW20, Definition 7.18]). In these quantum hypothesis testing tasks, one assumes that there are many copies of the states available and furthermore that it is possible to perform a collective measurement on them. The technological capabilities required to perform such a collective measurement appear to be quite challenging, and it seems one might generally need a quantum computer to do so [DSM+22, CDCS24].
The measured relative entropy [Don86, Pia09] is a distinguishability measure that relaxes the requirements of quantum hypothesis testing signficantly, and it has been generalized to the Rényi family as well [Fuc96, Eqs. (3.116)–(3.117)]. Indeed, the idea behind these measures is to evaluate classical distinguishability measures on the distributions that result from performing a measurement on a single copy of the state and then optimize them over all possible measurements. See Definition 1 and Definition 2 for precise definitions of the standard and Rényi measured relative entropies, respectively. In such a way, these quantities lead to technologically feasible strategies for quantum hypothesis testing, which consist of a hybrid approach involving quantum measurement and classical post-processing. Indeed, even though there are gaps between the fundamental error rates of quantum hypothesis testing under general, collective measurements and those that result from these hybrid quantum–classical strategies, the latter strategies are more feasible in the near term.
@@ 1.2 Summary of results
Indeed, for quantum states, our main contributions are to establish reductions of the variational formulas of [BFT17] to semi-definite optimization problems involving linear objective functions and the aforementioned hypographs or epigraphs (see Propositions 3 and 7). This finding is admittedly a rather direct combination of the contributions of [BFT17, FS17, FSP19]. However, it is ultimately useful in establishing our next contribution, which is an extension of these findings to measured relative entropies of channels. To establish these latter results, we use basic properties of weighted geometric means and the operator connection of the logarithm (see Propositions 10 and 12).
One benefit of our findings is that they lead to semi-definite programs involving linear matrix inequalities each of size $2d\times 2d$ when the states are $d\times d$ matrices and of size $2d_{A}d_{B}\times 2d_{A}d_{B}$ when the channels have input dimension $d_{A}$ and output dimension $d_{B}$ (see Propositions 3, 7, 10, and 12 for precise statements). As such, they do not suffer from the quadratic increase in size that occurs when applying the approach from [FS17, FSP19] to the Petz–Rényi and standard quantum relative entropies (however, note that there has been progress on addressing this issue more recently [FS23, FF23, KS24]). Furthermore, it is unclear how to apply the approach from [FS17, FSP19] for computing the dynamical (channel) version of these quantities. However, one of our main contributions is semi-definite programs for the measured relative entropies of channels, and the transition from our claims for states to our claims for channels is smooth, with the proofs consisting of just a few lines (see (71)–(80) and (89)–(98) for these steps).
@@ 1.3 Organization of the paper
The rest of our paper is organized as follows. Section 2 establishes notation and reviews background material, including the weighted geometric mean and its properties, its hypograph and epigraph, and operator connections and their properties (especially for the logarithm). The remaining Sections 3, 4, 5, and 6 provide essential definitions and detail our main results for measured Rényi relative entropy of states, measured relative entropy of states, measured Rényi relative entropy of channels, and measured relative entropy of channels, respectively. We conclude in Section 7 with a brief summary and some directions for future research.
@@ 3.1 Definition and basic properties
For $\alpha>1$, the measured Rényi relative entropy is finite if and only if $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$. If the support condition holds, then it follows that the support of $\left(\operatorname{Tr}[\Lambda_{x}\rho]\right)_{x\in\mathcal{X}}$ is contained in the support of $\left(\operatorname{Tr}[\Lambda_{x}\sigma]\right)_{x\in\mathcal{X}}$, which in turn implies that $D_{\alpha}^{M}(\rho\|\sigma)<+\infty$. If the support condition does not hold, then $D_{\alpha}^{M}(\rho\|\sigma)=+\infty$, by applying the argument in [RSB24, Section 3.5].
We now recall some basic properties of the measured Rényi relative entropies, the first of which is actually a consequence of [BFT17, Theorem 4] and the second observed in [RSB24, Lemma 5].
@@ 4.1 Definition and basic properties
Let us note that the convergence statement above can be made more precise by invoking (22) and [CDL+25, Lemma 23] (see also [BDSW24, Lemma 13]). Specifically, when $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$, there exists a state-dependent constant $c(\rho\|\sigma)$ such that, for all $\delta\in(0,\frac{\ln 3}{2c(\rho\|\sigma)}]$, the following bound holds:
@@ 4.3 Optimizing the measured relative entropy of states
$D^{M}(\rho\|\sigma)=\sup_{\omega>0,\theta\in\operatorname{Herm}}\left\{\operatorname{Tr}[\theta\rho]-\operatorname{Tr}[\omega\sigma]+1:\theta\leq P_{\ln}(I,\omega)\right\},$ (59)
@@
Furthermore, when $\rho$ and $\sigma$ are $d\times d$ matrices, the quantity $D^{M}(\rho\|\sigma)$ can be efficiently calculated by means of a semi-definite program up to an additive error $\varepsilon$, by means of $O(\sqrt{\ln(1/\varepsilon)})$ linear matrix inequalities, each of size $2d\times 2d$.
Proof. The formula above is a direct consequence of (58) and the following identity:
@@
As such, we have rewritten $D^{M}(\rho\|\sigma)$ in terms of the hypograph of $P_{\ln}(I,\omega)$. By appealing to [FSP19, Theorem 2 and Theorem 3], it follows that $D^{M}(\rho\|\sigma)$ can be efficiently calculated by means of a semi-definite program with the stated complexity.
@@
An alternative approach for computing the measured relative entropy is to set $\alpha=1-2^{-\ell}$ for $\ell\in\mathbb{N}$, similar to what was done in [FF21a, Lemma 9]. By appealing to Proposition 3 and (54), it follows that, in order to achieve an error $\varepsilon$ in computing the measured relative entropy, the semi-definite program resulting from this approach requires $O(\ln(1/\varepsilon))$ linear matrix inequalities, each of size $2d\times 2d$. When compared to the performance of the approach from Proposition 7, it is clear that this latter approach is preferred because it requires only $O(\sqrt{\ln(1/\varepsilon)})$ linear matrix inequalities to achieve the same error.

### v3

@@ 1.1 Background
In many quantum information processing tasks, distinguishing quantum states or channels is central to assessing system performance. Measures of distinguishability such as the relative entropy [KL51] and its generalization to Rényi relative entropy [R6́1] play important roles in information theory, finding direct operational meaning in hypothesis testing tasks [Che52, Ste, Che56, Str62, Hoe65, HK89] while being used to construct other entropic measures like mutual information and conditional entropy [Sib69, Aug78, Csi95]. There are a number of quantum generalizations of these quantities [Ume62, Pet85, Pet86, MLDS+13, WWY14], finding operational meaning in quantum hypothesis testing tasks [HP91, ON00, Nag06, ACM+07, Hay07, ANSV08, NS09] while being used to construct other entropic measures like quantum mutual information and conditional entropy (see, e.g., [KW20, Definition 7.18]).
Typically, the optimal hypothesis-testing success rate associated with the quantum relative entropy (Umegaki relative entropy) requires collective measurements over many identical copies of a state. Implementing such joint measurements, however, demands substantial resources—potentially full-scale quantum computation [DSM+22, CDCS24]. In contrast, measured relative entropies [Don86, Pia09] significantly relax these technological demands. Intuitively, the ‘measured’ versions of quantum relative entropies quantify how well two quantum states can be distinguished by performing measurements on individual copies of the state, as opposed to using an arbitrary collective quantum measurement. Instead of collective operations, they quantify how well two states (or channels) can be distinguished when one performs a measurement on each system separately. In this way, the measured quantities capture the fundamental limits of hybrid quantum–classical discrimination strategies that are achievable with today’s experimental capabilities.
The measured relative entropy [Don86, Pia09] has been generalized to the Rényi family as well [Fuc96, Eqs. (3.116)–(3.117)]. Indeed, the idea behind these measures is to evaluate classical distinguishability measures on the distributions that result from performing a measurement on a single copy of the state and then optimize them over all possible measurements. See Definition 1 and Definition 2 for precise definitions of the standard and Rényi measured relative entropies, respectively. In such a way, these quantities lead to technologically feasible strategies for quantum hypothesis testing, which consist of a hybrid approach involving quantum measurement and classical post-processing. Indeed, even though there are gaps between the fundamental error rates of quantum hypothesis testing under general, collective measurements and those that result from these hybrid quantum–classical strategies, the latter strategies are more feasible in the near term.
@@ 1.2 Summary of results
Indeed, for quantum states, our main contributions are to establish reductions of the variational formulas of [BFT17] to semi-definite optimization problems involving linear objective functions and the aforementioned hypographs or epigraphs (see Propositions 3 and 7). This finding is admittedly a rather direct combination of the contributions of [BFT17, FS17, FSP19]. However, it is ultimately useful in establishing our next contribution, which is an extension of these findings to measured relative entropies of channels. To establish these latter results, we use basic properties of weighted geometric means and the operator connection of the logarithm (see Propositions 11 and 13).
One benefit of our findings is that they lead to semi-definite programs involving linear matrix inequalities each of size $2d\times 2d$ when the states are $d\times d$ matrices and of size $2d_{A}d_{B}\times 2d_{A}d_{B}$ when the channels have input dimension $d_{A}$ and output dimension $d_{B}$ (see Propositions 3, 7, 11, and 13 for precise statements). As such, they do not suffer from the quadratic increase in size that occurs when applying the approach from [FS17, FSP19] to the Petz–Rényi and standard quantum relative entropies (however, note that there has been progress on addressing this issue more recently [FS23, FF23, KS24]). Furthermore, it is unclear how to apply the approach from [FS17, FSP19] for computing the dynamical (channel) version of these quantities. However, one of our main contributions is semi-definite programs for the measured relative entropies of channels, and the transition from our claims for states to our claims for channels is smooth, with the proofs consisting of just a few lines (see (105)–(114) and (125)–(134) for these steps).
@@ 1.3 Organization of the paper
The rest of our paper is organized as follows. Section 2 establishes notation and reviews background material, including the weighted geometric mean and its properties, its hypograph and epigraph, and operator connections and their properties (especially for the logarithm). The remaining Sections 34, 5, and 6 provide essential definitions and detail our main results for measured Rényi relative entropy of states, measured relative entropy of states, measured Rényi relative entropy of channels, and measured relative entropy of channels, respectively. We conclude in Section 7 with a brief summary and some directions for future research.
@@ 3.1 Definition and basic properties
For $\alpha>1$, the measured Rényi relative entropy is finite if and only if $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$. If the support condition holds, then it follows that the support of $\left(\operatorname{Tr}[\Lambda_{x}\rho]\right)_{x\in\mathcal{X}}$ is contained in the support of $\left(\operatorname{Tr}[\Lambda_{x}\sigma]\right)_{x\in\mathcal{X}}$, which in turn implies that $D_{\alpha}^{M}(\rho\|\sigma)<+\infty$. If the support condition does not hold, then $D_{\alpha}^{M}(\rho\|\sigma)=+\infty$, by applying the argument in [RSB25, Section 3.5].
We now recall some basic properties of the measured Rényi relative entropies, the first of which is actually a consequence of [BFT17, Theorem 4] and the second observed in [RSB25, Lemma 5].
@@ 4.1 Definition and basic properties
Let us note that the convergence statement in Proposition 4 can be made more precise by invoking (22) and [CDL+25, Lemma 23] (see also [BDSW24, Lemma 13]). Specifically, when $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$, there exists a state-dependent constant $c(\rho\|\sigma)$ such that, for all $\delta\in(0,\frac{\ln 3}{2c(\rho\|\sigma)}]$, the following bound holds:
@@ 4.3 Optimizing the measured relative entropy of states
$D^{M}(\rho\|\sigma)=\sup_{\omega>0,\theta\in\mathbb{H}}\left\{\operatorname{Tr}[\theta\rho]-\operatorname{Tr}[\omega\sigma]+1:\theta\leq P_{\ln}(I,\omega)\right\},$ (59)
@@
Furthermore, when $\rho$ and $\sigma$ are $d\times d$ matrices, the quantity $D^{M}(\rho\|\sigma)$ can be efficiently calculated by means of a semi-definite program up to an additive error $\varepsilon$, by means of $O(\sqrt{\ln(1/\varepsilon)})$ linear matrix inequalities, each of size $2d\times 2d$. Specifically, there exist $m,k\in\mathbb{N}$ such that $m+k=O(\sqrt{\ln(1/\varepsilon)})$ and the following inequality holds:
$\left|D^{M}(\rho\|\sigma)-D^{M}_{m,k}(\rho\|\sigma)\right|\leq\varepsilon,$ (60)
where
$D_{m,k}^{M}(\rho\|\sigma)\coloneqq\sup_{\begin{subarray}{c}\omega>0,\theta\in\mathbb{H},\\ T_{1},\ldots,T_{m}\in\mathbb{H},\\ Z_{0},\ldots,Z_{k}\in\mathbb{H}\end{subarray}}\left\{\begin{array}[]{c}\operatorname{Tr}\!\left[\theta\rho\right]-\operatorname{Tr}\!\left[\omega\sigma\right]+1:\\ Z_{0}=\omega,\,\sum_{j=1}^{m}w_{j}T_{j}=2^{-k}\theta,\\ \left\{\begin{bmatrix}Z_{i}&Z_{i+1}\\ Z_{i+1}&I\end{bmatrix}\geq 0\right\}_{i=0}^{k-1},\\ \left\{\begin{bmatrix}Z_{k}-I-T_{j}&-\sqrt{t_{j}}T_{j}\\ -\sqrt{t_{j}}T_{j}&I-t_{j}T_{j}\end{bmatrix}\geq 0\right\}_{j=1}^{m}\end{array}\right\}$ (61)
and, for all $j\in\left\{1,\ldots,m\right\}$, $w_{j}$ and $t_{j}$ are the weights and nodes, respectively, for the $m$-point Gauss–Legendre quadrature on the interval $\left[0,1\right]$.
Proof. The formula in (59) is a direct consequence of (58) and the following identity:
@@
As such, we have rewritten $D^{M}(\rho\|\sigma)$ in terms of the hypograph of $P_{\ln}(I,\omega)$. By appealing to [FSP19, Theorem 2 and Theorem 3], it follows that $D^{M}(\rho\|\sigma)$ can be efficiently calculated by means of a semi-definite program with the stated complexity. In more detail, following [FSP19, Eqs. (9) and (11)], let us define the functions $r_{m}(x)$ and $r_{m,k}(x)$ as follows:
$r_{m}(x)$ $\coloneqq\sum_{j=1}^{m}w_{j}\frac{x-1}{t_{j}\left(x-1\right)+1},$ (63)
$r_{m,k}(x)$ $\coloneqq 2^{k}r_{m}(x^{1/2^{k}}),$ (64)
where $w_{j}$ and $t_{j}$ are as stated above. We can then define the following perspective function:
$P_{r_{m,k}}(X,Y)\coloneqq X^{1/2}r_{m,k}(X^{-1/2}YX^{-1/2})X^{1/2},$ (65)
which is jointly concave in $X$ and $Y$ [FSP19, Proposition 3], and its hypograph cone as follows:
$\operatorname{hyp}_{r_{m,k}}\coloneqq\left\{\left(X,Y,T\right)\in\mathbb{P}_{>0}(\mathcal{H})\times\mathbb{P}_{>0}(\mathcal{H})\times\mathbb{H}(\mathcal{H}):P_{r_{m,k}}(X,Y)\geq T\right\}.$ (66)
Then [FSP19, Theorem 3] states that $\left(X,Y,T\right)\in\operatorname{hyp}_{r_{m,k}}$ if and only if there exist $T_{1},\ldots,T_{m},Z_{0},\ldots,Z_{k}\in\mathbb{H}$ such that
$Z_{0}$ $=Y,$ (67)
$\sum_{j=1}^{m}w_{j}T_{j}$ $=2^{-k}T,$ (68)
$\begin{bmatrix}Z_{i}&Z_{i+1}\\ Z_{i+1}&X\end{bmatrix}$ $\geq 0\qquad\forall i\in\left\{0,\ldots,k-1\right\},$ (69)
$\begin{bmatrix}Z_{k}-X-T_{j}&-\sqrt{t_{j}}T_{j}\\ -\sqrt{t_{j}}T_{j}&X-t_{j}T_{j}\end{bmatrix}$ $\geq 0\qquad\forall j\in\left\{1,\ldots,m\right\}.$ (70)
As such, the quantity $D_{m,k}^{M}(\rho\|\sigma)$, as defined in (61), is equal to
$D_{m,k}^{M}(\rho\|\sigma)=\sup_{\omega>0,\theta\in\mathbb{H}}\left\{\operatorname{Tr}\!\left[\theta\rho\right]-\operatorname{Tr}\!\left[\omega\sigma\right]+1:\theta\leq P_{r_{m,k}}(I,\omega)\right\}.$ (71)
To arrive at the claim in (60), we invoke Lemma 8, which is a straightforward matrix generalization of [FSP19, Theorem 2]. Specifically, by invoking the first part of Lemma 8, we conclude that
$\theta\leq P_{\ln}(I,\omega)\quad\implies\quad\theta-\varepsilon I\leq P_{r_{m,k}}(I,\omega),$ (72)
so that
$D^{M}(\rho\|\sigma)$ $=\sup_{\omega>0,\theta\in\mathbb{H}}\left\{\operatorname{Tr}\!\left[\theta\rho\right]-\operatorname{Tr}\!\left[\omega\sigma\right]+1:\theta\leq P_{\ln}(I,\omega)\right\}$ (73)
$\leq\sup_{\omega>0,\theta\in\mathbb{H}}\left\{\operatorname{Tr}\!\left[\theta\rho\right]-\operatorname{Tr}\!\left[\omega\sigma\right]+1:\theta-\varepsilon I\leq P_{r_{m,k}}(I,\omega)\right\}$ (74)
$=\sup_{\omega>0,\theta^{\prime}\in\mathbb{H}}\left\{\operatorname{Tr}\!\left[\left(\theta^{\prime}+\varepsilon I\right)\rho\right]-\operatorname{Tr}\!\left[\omega\sigma\right]+1:\theta^{\prime}\leq P_{r_{m,k}}(I,\omega)\right\}$ (75)
$=\sup_{\omega>0,\theta^{\prime}\in\mathbb{H}}\left\{\operatorname{Tr}\!\left[\theta^{\prime}\rho\right]-\operatorname{Tr}\!\left[\omega\sigma\right]+1:\theta^{\prime}\leq P_{r_{m,k}}(I,\omega)\right\}+\varepsilon$ (76)
$=D_{m,k}^{M}(\rho\|\sigma)+\varepsilon.$ (77)
The second equality follows from the substitution $\theta^{\prime}=\theta-\varepsilon I$ and the penultimate equality follows because $\operatorname{Tr}\!\left[\left(\theta^{\prime}+\varepsilon I\right)\rho\right]=\operatorname{Tr}\!\left[\theta^{\prime}\rho\right]+\varepsilon$. Now by invoking the second part of Lemma 8, we conclude that
$\theta\leq P_{r_{m,k}}(I,\omega)\quad\implies\quad\theta-\varepsilon I\leq P_{\ln}(I,\omega),$ (78)
so that
$D_{m,k}^{M}(\rho\|\sigma)$ $=\sup_{\omega>0,\theta\in\mathbb{H}}\left\{\operatorname{Tr}\!\left[\theta\rho\right]-\operatorname{Tr}\!\left[\omega\sigma\right]+1:\theta\leq P_{r_{m,k}}(I,\omega)\right\}$ (79)
$\leq\sup_{\omega>0,\theta\in\mathbb{H}}\left\{\operatorname{Tr}\!\left[\theta\rho\right]-\operatorname{Tr}\!\left[\omega\sigma\right]+1:\theta-\varepsilon I\leq P_{\ln}(I,\omega)\right\}$ (80)
$=\sup_{\omega>0,\theta^{\prime}\in\mathbb{H}}\left\{\operatorname{Tr}\!\left[\left(\theta^{\prime}+\varepsilon I\right)\rho\right]-\operatorname{Tr}\!\left[\omega\sigma\right]+1:\theta^{\prime}\leq P_{\ln}(I,\omega)\right\}$ (81)
$=\sup_{\omega>0,\theta^{\prime}\in\mathbb{H}}\left\{\operatorname{Tr}\!\left[\theta^{\prime}\rho\right]-\operatorname{Tr}\!\left[\omega\sigma\right]+1:\theta^{\prime}\leq P_{\ln}(I,\omega)\right\}+\varepsilon$ (82)
$=D^{M}(\rho\|\sigma)+\varepsilon.$ (83)
Combining (77) and (83), we conclude (60).
Lemma 8
Let $a>1$ and $\varepsilon>0$. Then there exist $m,k\in\mathbb{N}$ with $m+k=O(\sqrt{\ln(1/\varepsilon)})$ such that
  1. 1.
    If $0\leq a^{-1}Y\leq X\leq aY$ and $\left(X,Y,T\right)\in\operatorname{hyp}_{\ln}$, then $\left(X,Y,T-X\varepsilon\right)\in\operatorname{hyp}_{r_{m,k}}.$
  2. 2.
    If $0\leq a^{-1}Y\leq X\leq aY$ and $\left(X,Y,T\right)\in\operatorname{hyp}_{r_{m,k}}$, then $\left(X,Y,T-X\varepsilon\right)\in\operatorname{hyp}_{\ln}.$
Proof. The proof follows the approach given in the proof of [FSP19, Theorem 6]. Consider that
$0$ $\leq a^{-1}Y\leq X\leq aY$ (84)
$\implies\qquad 0$ $\leq a^{-1}I\leq Y^{-1/2}XY^{-1/2}\leq aI$ (85)
$\implies\qquad 0$ $\leq a^{-1}I\leq Y^{1/2}X^{-1}Y^{1/2}\leq aI$ (86)
$\implies\qquad 0$ $\leq a^{-1}I\leq X^{-1/2}YX^{-1/2}\leq aI,$ (87)
where (85) follows by left and right multiplying (84) by $Y^{-1/2}$, (86) follows from operator antimonotonicity of the function $x\mapsto x^{-1}$, specifically,
$a^{-1}I\leq Y^{-1/2}XY^{-1/2}$ $\implies Y^{1/2}X^{-1}Y^{1/2}\leq aI,$ (88)
$Y^{-1/2}XY^{-1/2}\leq aI$ $\implies a^{-1}I\leq Y^{1/2}X^{-1}Y^{1/2},$ (89)
and (87) follows because $Y^{1/2}X^{-1}Y^{1/2}$ and $X^{-1/2}YX^{-1/2}$ have the same spectrum. By [FSP19, Theorem 1], there exist $m,k\in\mathbb{N}$ with $m+k=O(\sqrt{\ln(1/\varepsilon)})$ such that
$\left|r_{m,k}(z)-\ln(z)\right|\leq\varepsilon,$ (90)
for all $z\in\left[a^{-1},a\right]$. Then applying this to (87) gives that
$r_{m,k}(X^{-1/2}YX^{-1/2})-\ln(X^{-1/2}YX^{-1/2})\geq-\varepsilon I,$ (91)
which implies that
$X^{1/2}r_{m,k}(X^{-1/2}YX^{-1/2})X^{1/2}-X^{1/2}\ln(X^{-1/2}YX^{-1/2})X^{1/2}\geq-\varepsilon X.$ (92)
By assumption, $\left(X,Y,T\right)\in\operatorname{hyp}_{\ln}$, so that
$X^{1/2}\ln(X^{-1/2}YX^{-1/2})X^{1/2}\geq T.$ (93)
Now adding (92) and (93) gives
$X^{1/2}r_{m,k}(X^{-1/2}YX^{-1/2})X^{1/2}\geq T-\varepsilon X,$ (94)
which is equivalent to $P_{r_{m,k}}(X,Y)\geq T-\varepsilon X$, thus establishing the first claim. The proof of the second claim is similar.
@@
An alternative approach for computing the measured relative entropy is to set $\alpha=1-2^{-\ell}$ for $\ell\in\mathbb{N}$, similar to what was done in [FF21a, Lemma 9]. By appealing to Proposition 3 and (54), it follows that, in order to achieve an additive error $\varepsilon$ in computing the measured relative entropy, the semi-definite program resulting from this approach requires $O(\ln(1/\varepsilon))$ linear matrix inequalities, each of size $2d\times 2d$. When compared to the performance of the approach from Proposition 7, it is clear that this latter approach is preferred because it requires only $O(\sqrt{\ln(1/\varepsilon)})$ linear matrix inequalities in order to achieve the same error.